Derive the expression for calculating the wavelength () in terms of y1bright. Given the ratio of d to D is 1:1000 and also the angle 1 is 7
Answer #1

Since the angle is 7deg we can use the small-angle approximation

   m?/d = y/D

. Since the y value is called y1bright, we assume we are looking at the 1st interference maximum, so m = 1.

Then ? = yd/D

          = y(1/1000)

          = 0.001y